# Vibe coding: programming through conversation with artificial intelligence

### v1

@@ 3.4.1 How do creators set up a vibe coding session for effective prompting?
The creator in TW1 takes a formalised approach to prompt customisation, starting with an existing template called .cursorrules sourced from GitHub. They incorporate this into their own Claude.md file, which serves as system instructions for the code-generating model. They personalise these instructions by specifying the code languages, frameworks, and libraries used in their codebase.
@@ 3.4.2 What explicit prompting strategies are used during vibe coding?
As mentioned, the developer in YT22 began with documentation and implementation plans generated by Code Guide as context for prompting Cursor. TW1 also demonstrates a meta-prompting approach by using one model to generate a prompt for the coding AI agent. The developer in YT22 also adapts community knowledge, using a prompt structure found on Twitter but tailoring it to his specific project context. He explicitly practices scope limitation, focusing on one “phase” of the implementation plan at a time to reduce hallucinations. In another form of scope limitation, he also introduces constraints to prompts such as “don’t integrate stripe yet. Just make a design with dummy data”, perhaps anticipating and preempting the model’s tendency to fill in unstated gaps, or tendency to set an overambitious goal that the model is likely to fail to achieve in a single step.
The developer in YT22 demonstrates further fine-grained guidance methods, particularly by providing external documentation as context, explicitly copying the URL of the official Tanstack Form documentation with instructions to “use docs,” employing the model’s ability to index external resources. When encountering runtime issues, he supplements prompts with error messages, directly pasting browser errors and asking for fixes.
@@
The creator in YT15 uses both precise and imprecise language to convey aesthetics. When requesting arrow repositioning, the creator gives specific directions like “not to use diagonal ones” and that “arrowheads should be pointing to the middle of the boxes.” However, the creator also employs imprecise descriptors such as expanding space “significantly,” relying on the AI’s interpretation rather than providing exact measurements. The creator also issues a high-level prompt mentioning a theme and colours without specifying details, leaving these to the “interpretation” of the model. Similarly, the programmer in YT22 uses a mix of specific and ambiguous instructions for styling, such as “move the date below and use a very small font” and “don’t use bold font, narrow the space between conversations.” These prompts combine specific actions with somewhat subjective descriptors like “very small” and “narrow,” leaving some interpretation to the model while maintaining directional control.
YT21 shows a similar pattern but with explicit reflection on the strategy. The programmer initially uses very brief, high-level prompts such as simply requesting to “create an API to create conversation.” When these broad prompts fail to yield the desired results, the programmer acknowledges, “I should have been more clear here,” and shifts to specific, low-level prompts that target particular functions, variables, and styling elements. The programmer also employs extremely fine-grained iterations for localised changes, such as using inline prompts to modify specific parts of functions. The programmer’s statement that he “starts by asking very dumb questions. And if it gets confused, [he] starts providing more detail” reveals an intentional approach to granularity adjustment.
The developer in TW1 attempts to keep prompting succinct yet detailed, but ultimately finds longer prompts to be better so as not to give the AI “a lot of leeway to go off the rails.” This developer strikes a balance between high-level directives (e.g, “Increase the exp for VIP members to 1.3x on the back-end”) while providing specific implementation details when necessary (e.g., which files to edit, and specific callouts about variables or code types within these broader requests).
YT22 intentionally modifies high-level prompts to focus on smaller chunks (“just do phase one”) to reduce hallucinations, revealing a mental model of how prompt granularity affects AI performance. Their debugging prompts were consistently single-objective (pasting specific errors with requests to fix), and even feature requests focused on particular elements like "Make a page with all resumes that run through the optimization" rather than requesting multiple features simultaneously, demonstrating a preference for manageable scope and sequential problem-solving. YT22 also reveals situations where extremely fine-grained prompting may be deemed inefficient; the creator often makes manual edits rather than prompting for minor adjustments.
@@ 3.4.4 What is the role of different input modes such as voice, text, and images, in vibe coding prompting?
Code reuse strategies appear consistently across the examined videos, naturally through textual means. In YT21, the creator frequently copied and pasted entire function bodies into the composer window along with text instructions to provide context. Similarly, the creator would copy error messages from the terminal or browser console into the prompt window. The inclusion of expected output formats also appears as a code reuse strategy in YT21, where the creator provided an example of the expected JSON response to guide the model.
@@ 3.5.1 What debugging strategies are applied during vibe coding?
More technically sophisticated debugging approaches appear in YT21 and YT22, where the programmers heavily utilized browser developer tools. In YT21, the programmer examined error messages in the browser’s console and network tabs to pinpoint issues and verify API calls. Similarly, in YT22, when encountering a blank screen, the creator immediately opened the browser’s developer console, identifying a “trpc error”. Terminal analysis was also observed in YT21, with the programmer monitoring terminal output alongside browser tools.
@@
Creators frequently employ iterative refinement through targeted prompting. In YT15, the creator provided detailed instructions for repositioning arrows, fixing overlapping labels, and adjusting animations. This strategy of using natural language prompts to guide debugging is also observed in YT22, where instead of directly modifying code, the creator issued prompts such as “I don’t need the function of pasting the URL of job description. Please use the text approach only.” TW1 employed a strategy of resubmitting a prompt after iteration on the system instructions to guide the model to produce results inline with her preferences.
Despite the AI assistance, creators still engage in manual code review and editing. In YT21, the programmer frequently scanned and read the AI-generated code to assess alignment with expectations and identify potential issues before testing. Similarly, in YT22, the creator frequently examines the generated code, scrolling through files and looking at diffs to understand the logic of the AI output. In TW1, the creator inspected the AI-generated code to check if it was viable, and in some cases if the output was “close”, they accepted the code with the intention to fix it manually.
@@
The special nature of code reading in vibe coding
@@
Experienced programmers can look at the AI-generated code and intuitively sense if it’s using the right level of abstraction. Observing the AI writing low-level HTML/JS/CSS primitives for diagrams, the programmer immediately recognized the pattern and felt like they were “reinventing the wheel”, speculating there “should be some libraries that do these things” (YT15).
@@ 3.6.1 What challenges besides code bugs do programmers encounter while vibe coding?
The observations reveal several challenges that practitioners face when engaging in vibe coding beyond traditional code bugs. These obstacles span technical, conceptual, and workflow dimensions.
One challenge is communicating visual and abstract ideas. In YT15, the creator struggled with articulating what he wanted to do with the arrows, encountering the fundamental difficulty of trying to talk about diagrams using words. This communication barrier extends to complex conceptual goals as well, as seen in TW1 where the developer faced difficulties forming the prompt to match what she wanted for a matchmaking feature, asking “how is [AI] going to give me what I want if I don’t even know?”
Understanding AI capabilities and limitations represents another challenge. In YT15, the creator expressed uncertainty about whether a specific AI model (Claude 3.5 Sonnet) could create an animation as desired and later switched to a different model (o3 mini) after an unsatisfactory experience. In TW1, the creator felt that some models were better at certain workflows, indicating that matching the right AI capability to the right task requires experience and knowledge with the models. TW1 also stated that “LLMs will get you most of the way there, but they don’t get you all the way there”, perceiving universally inherent limitations in AI capabilities.
AI hallucinations and failures to fully adhere to prompt instructions create obstacles. In YT21, the model sometimes generated code with non-existent properties, requiring the developer to identify these issues through error messages and code inspection. The developer in TW1 expressed frustration when generated code did not adhere to the system instructions she defined, noting it exported a function when it didn’t need to. In YT22, the model produced documentation referencing Next.js 14 and JavaScript files when the creator was using Next.js 15 with TypeScript, requiring manual updates. The need to include specific documentation for Tanstack Form to ensure proper integration also demonstrated that the model’s “general knowledge” isn’t always sufficient for specialised frameworks.
@@
Deciding when to rely on AI versus manual intervention introduces new metacognitive load (Tankelevitch et al.,, 2024). In YT21, the programmer frequently reviewed and sometimes rejected AI-generated code, stating that reviewing changes helped him remain in control. In YT22, the creator frequently switched between prompting and making manual edits without clear criteria for choosing one approach over the other. The developer in TW1 observed that for small adjustments, she felt faster working manually, but for more substantial edits, she preferred using AI assistance.
@@ 3.6.2 How are these challenges addressed?
Iterative refinement of prompts is inevitably the programmers’ first recourse when encountering communication challenges with the model. In YT15, when the creator struggled to articulate visual intent for arrow placement, they iterated through multiple prompts with increasingly detailed spoken instructions. Similarly, in YT21, the programmer recognised when initial prompts were too vague and provided more detailed and specific prompts in subsequent attempts, often including numbered lists of requirements and referencing existing code snippets. When the model “did not listen to the system instructions”, TW1 refined the system instructions to emphasise acceptable behaviour; a higher-order strategy than refining individual prompts.
At one point, the programmer in TW1 adopted an unusual prompting strategy where the developer responded to an unsatisfactory result by prompting “Bro, you are losing aura with this. Come on.” before pasting the original prompt and resubmitting. This explicit “vibification” produced more code than the previous turn, though the programmer in TW1 still expressed dissatisfaction with the code quality. This suggests that some vibe coders experiment with social or emotional appeals, albeit perhaps tongue-in-cheek, in prompts. We also observed similar styles of prompting in other videos in our broader corpus. The results of these, as might be expected, are mixed.
Developers strategically switch between different AI models based on their perceived strengths. In YT15, the creator actively switched between OpenAI’s o3-mini and Anthropic’s Claude 3.5 Sonnet models based on his qualitative assessment and the perceived strengths of each for different aspects of the task, using one for animation and another for initial code generation. Similarly, in TW1, the developer discussed where she used models from the Claude family to generate code for a web page, but switched to OpenAI’s GPT-4.5 to rewrite the text because she believed its text quality was an improvement over Claude.
@@ 3.7.1 When and in what ways do experts deploy their expertise in vibe coding?
Expert knowledge also guides problem identification and solution development. In YT22, upon encountering a blank screen, the creator quickly inspects the browser console and identifies a ‘trpc error,’ demonstrating expertise in reading and interpreting error messages. In YT21, the programmer displays expert knowledge of debugging techniques by immediately recognising database issues from error messages, effectively using browser developer tools to diagnose problems, forming hypotheses about root causes, and inserting print statements to trace code execution. In YT21, the programmer demonstrates understanding of API endpoints, data structures, and how different parts of the application communicate, allowing them to identify issues with API calls. Similarly, in YT22, the creator identifies that an application is trying to connect to PostgreSQL instead of Supabase and knows exactly how to manually edit environment files to correct the configuration. In YT22, the creator evaluates AI-generated code against expected standards by checking generated React components against documentation, demonstrating and applying their knowledge of UI libraries.
Expertise guides quality control and feature alignment. In YT22, the creator recognizes and decides to remove unwanted AI-generated features that, while not buggy or intrusive, don’t align with the intended design. In TW1, the creator demonstrates understanding of when generated code will add tech debt which “lowers developer velocity”, and addresses these issues proactively.
@@ 3.7.2 What kinds of expertise are used in vibe coding?
Beyond “traditional” programming expertise (code quality and maintainability, fault localisation and debugging), the programmers we studied exhibited expertise in two additional domains: AI expertise (understanding models, prompting practices, concepts such as context windows, and limitations of AI code generation), and product management expertise (forming and translating goals into features for the product). Perhaps as a consequence of the blending of prototyping and production (as discussed in Section 3.2.2), vibe coding requires (or at least is greatly facilitated by) a multifaceted skillset spanning technical implementation, tool understanding, and product vision.
@@
Moreover, we speculate that these programmers are also developing a sense of what might be called ambient competence. This is a sense of competence wherein the programmer feels capable of tackling tasks they wouldn’t have considered before, because they have access to an AI system that may be able to accomplish the task with very little effort invested, regardless of whether they themselves have the skills to tackle them manually (which they very well may). This can be thought of as the competence that arises as the result of the active fulfilment of the “awareness of the possible” (Sarkar, 2023b, ; Sarkar, 2023a, ). This could lead to a shifting locus of agency, where the programmer’s confidence and the scope and complexity of projects they undertake are increasingly defined by their perceived ability to prompt and guide a model, rather than their own direct coding abilities.
@@ 3.7.3 When do practitioners transition from vibe coding to manual work?
For straightforward edits where the overhead of prompting exceeds the effort of making the change directly, developers often opt for manual work. In YT21, the programmer attempted an inline edit with AI but then decided to perform the simple one-line edit manually after determining that AI assistance was unnecessary for such a minor change. In TW1, the developer articulates her decision-making process about when to work manually versus using AI assistance. In one instance, she switches to manual work when she believed it would be a more efficient workflow, noting that “IntelliJ’s autocomplete and search is still faster than LLMs”. This leads her to “race the AI” to find code locations, sometimes stopping AI generation mid-process when she locates what she needs faster.
@@
Another common transition point across videos is when AI generates a mostly-complete solution that requires refinement. This can be viewed as a special case of transitioning to manual work for efficiency. For example, TW1 accepted code when the AI “got close” to her goal and then manually edited the generated code to be more inline with her preferences for code. Similarly, in YT22, the creator performs manual edits to remove unwanted UI elements, delete redundant buttons, and make small edits in the code to better align with their vision. TW1 also describes her viewpoint that LLMs should be used to “get close to what you want” but that expecting perfect outputs is “a mistake” – instead, the optimal workflow for her involved making post-generation adjustments to generated code, and reuse of generated code in similar features by quickly modifying code to other contexts rather than reprompt the AI. In YT22, the creator manually removes functionality for uploading job descriptions via URL to focus solely on text-based inputs, representing a conscious design decision that required direct code manipulation.
This refinement extends to project plans and documentation as well. In YT22, the developer manually corrects a misalignment in AI-generated documentation (changing "Next.js 14" to "Next.js 15") and deletes redundant steps in implementation plans.
@@
Debugging is as a common trigger for manual intervention across multiple videos. In YT21, the programmer switches to manual work almost immediately when encountering errors, using traditional debugging techniques like browser developer tools, console inspection, and hypothesis formation. In one instance, they hypothesised a solution based on an error message and implemented a fix without AI assistance (YT21). YT22 shows similar patterns, with the developer manually editing environment files to fix database connection errors and directly intervening to correct runtime issues. These examples suggest that when errors occur, developers often rely on their expertise and traditional debugging approaches rather than AI tools.
@@
However, YT15 presents a contrasting case where the developer states explicitly that he’s “not going to try and fix it on his own” when encountering an error, citing adherence to the Karpathy canon. Thus, the videos do exhibit individual differences in approaches to manual work. The creator in YT15 shows a deliberate avoidance of manual debugging and relies almost exclusively on re-prompting for changes throughout their workflow. In contrast, the programmer in YT21 frequently engages in manual code review and rejection of AI suggestions based on their understanding of correctness and project requirements.

### v2

@@ 3.4.1 How do creators set up a vibe coding session for effective prompting?
The creator in TW1 takes a formalised approach to prompt customisation, starting with an existing template called .cursorrules sourced from GitHub. They incorporate this into their own Claude.md file, which serves as system instructions for the code-generating model. They personalise these instructions by specifying the programming languages, frameworks, and libraries used in their codebase.
@@ 3.4.2 What explicit prompting strategies are used during vibe coding?
As mentioned, the developer in YT22 began with documentation and implementation plans generated by Code Guide as context for prompting Cursor. TW1 also demonstrates a meta-prompting approach by using one model to generate a prompt for the coding AI agent. The developer in YT22 also adapts community knowledge, using a prompt structure found on Twitter but tailoring it to their specific project context. They explicitly practice scope limitation, focusing on one “phase” of the implementation plan at a time to reduce hallucinations. In another form of scope limitation, they also introduce constraints to prompts such as “don’t integrate stripe yet. Just make a design with dummy data”, perhaps anticipating and preempting the model’s tendency to fill in unstated gaps, or tendency to set an overambitious goal that the model is likely to fail to achieve in a single step.
The developer in YT22 demonstrates further fine-grained guidance methods, particularly by providing external documentation as context, explicitly copying the URL of the official Tanstack Form documentation with instructions to “use docs,” employing the model’s ability to index external resources. When encountering runtime issues, they supplement prompts with error messages, directly pasting browser errors and asking for fixes.
@@
The creator in YT15 uses both precise and imprecise language to convey aesthetics. When requesting arrow repositioning, the creator gives specific directions like “not to use diagonal ones” and that “arrowheads should be pointing to the middle of the boxes.” However, the creator also employs imprecise descriptors such as expanding space “significantly,” relying on the model’s interpretation rather than providing exact measurements. The creator also issues a high-level prompt mentioning a desired theme and colour palette without specifying details, leaving these to the interpretation of the model. Similarly, the programmer in YT22 uses a mix of specific and ambiguous instructions for styling, such as “move the date below and use a very small font” and “don’t use bold font, narrow the space between conversations.” These prompts combine specific actions with somewhat subjective descriptors like “very small” and “narrow,” leaving some interpretation to the model while maintaining directional control.
YT21 shows a similar pattern but with explicit reflection on the strategy. The programmer initially uses very brief, high-level prompts such as simply requesting to “create an API to create conversation.” When these broad prompts fail to yield the desired results, the programmer acknowledges, “I should have been more clear here,” and shifts to specific, low-level prompts that target particular functions, variables, and styling elements. The programmer also employs extremely fine-grained iterations for localised changes, such as using inline prompts to modify specific parts of functions. The programmer’s statement that they “start by asking very dumb questions. And if it gets confused, [they] start providing more detail” reveals an intentional approach to granularity adjustment.
The developer in TW1 attempts to keep prompting succinct yet detailed, but ultimately finds longer prompts to be better so as not to give the model “a lot of leeway to go off the rails.” This developer strikes a balance between high-level directives (e.g, “Increase the exp for VIP members to 1.3x on the back-end”) while providing specific implementation details when necessary (e.g., which files to edit, and specific callouts about variables or code types within these broader requests).
YT22 intentionally modifies high-level prompts to focus on smaller chunks (“just do phase one”) to reduce hallucinations, revealing a mental model of how prompt granularity affects AI performance. Their debugging prompts were consistently single-objective (pasting specific errors with requests to fix), and even feature requests focused on particular elements like “Make a page with all resumes that run through the optimization” rather than requesting multiple features simultaneously. YT22 also reveals situations where extremely fine-grained prompting may be deemed inefficient; the creator often makes manual edits rather than prompting for minor adjustments.
@@ 3.4.4 What is the role of different input modes such as voice, text, and images, in vibe coding prompting?
Code reuse strategies appear consistently across the examined videos, naturally through textual means. In YT21, the creator frequently copied and pasted entire function bodies into the composer window to provide context for their natural language instructions. Similarly, the creator would copy error messages from the terminal or browser console into the prompt window. The inclusion of expected output formats also appears as a code reuse strategy in YT21, where the creator provided an example of the expected JSON response to guide the model.
@@ 3.5.1 What debugging strategies are applied during vibe coding?
AI “hallucinations” and failures to fully adhere to prompt instructions can create errors. In YT21, the model sometimes generated code with non-existent properties, requiring the developer to identify these issues through error messages and code inspection. The developer in TW1 expressed frustration when generated code did not adhere to the system instructions they defined, noting that it exported a function when it didn’t need to. In YT22, the model produced documentation referencing Next.js 14 and JavaScript files when the creator was using Next.js 15 with TypeScript, requiring manual updates.
@@
More technically sophisticated debugging approaches appear in YT21 and YT22, where the programmers heavily utilised browser developer tools. In YT21, the programmer examined error messages in the browser’s console and network tabs to pinpoint issues and verify API calls. Similarly, in YT22, when encountering a blank screen, the creator immediately opened the browser’s developer console, identifying a “trpc error”. Terminal analysis was also observed in YT21, with the programmer monitoring terminal output alongside browser tools.
@@
Creators frequently employ iterative refinement through targeted prompting. In YT15, the creator provided detailed instructions for repositioning arrows, fixing overlapping labels, and adjusting animations. This strategy of using natural language prompts to guide debugging is also observed in YT22, where instead of directly modifying code, the creator issued prompts such as “I don’t need the function of pasting the URL of job description. Please use the text approach only.” TW1 employed a strategy of resubmitting a prompt after iteration on the system instructions to guide the model to produce results in line with their preferences.
Creators still engage in manual code review and editing. In YT21, the programmer frequently scanned and read the AI-generated code to assess alignment with expectations and identify potential issues before testing. Similarly, in YT22, the creator frequently examines the generated code, scrolling through files and looking at diffs to understand the logic of the AI output. In TW1, the creator inspected the AI-generated code to check if it was viable, and in some cases if the output was “close”, they accepted the code with the intention to fix it manually.
@@
The special nature of code reading in vibe coding.
@@
Experienced programmers can look at AI-generated code and intuitively sense if it’s using the right level of abstraction. Observing the AI writing low-level HTML/JS/CSS primitives for diagrams, the programmer immediately recognised the pattern and felt like they were “reinventing the wheel”, speculating that there “should be some libraries that do these things” (YT15).
@@ 3.6.1 What challenges besides code bugs do programmers encounter while vibe coding?
Practitioners face several challenges when engaging in vibe coding beyond traditional code bugs. These challenges span technical, conceptual, and workflow dimensions.
One challenge is communicating visual and abstract ideas. In YT15, the creator struggled with articulating what they wanted to do with the arrows, encountering the fundamental difficulty of trying to talk about diagrams using words. Complex conceptual goals can also be difficult to articulate as prompts, as seen in TW1 where the developer faced difficulties forming the prompt to describe a matchmaking feature, asking “how is [AI] going to give me what I want if I don’t even know?”
Understanding AI capabilities and limitations is another challenge. In YT15, the creator expressed uncertainty about whether a specific AI model (Claude 3.5 Sonnet) could create an animation as desired and later switched to a different model (o3 mini) after an unsatisfactory experience. In TW1, the creator felt that some models were better at certain workflows, indicating that matching the right AI capability to the right task requires experience and knowledge with the models. TW1 also stated that “LLMs will get you most of the way there, but they don’t get you all the way there”, perceiving universally inherent limitations in AI capabilities. YT22’s need to include specific documentation for Tanstack Form to ensure proper integration also demonstrated that the model’s “general knowledge” isn’t always sufficient for specialised frameworks.
@@
Deciding when to rely on AI versus manual intervention (discussed further in Section 3.7.3) introduces new metacognitive load (Tankelevitch et al.,, 2024). In YT21, the programmer frequently reviewed and sometimes rejected AI-generated code, stating that reviewing changes helped them remain in control. In YT22, the creator frequently switched between prompting and making manual edits without clear criteria for choosing one approach over the other. The developer in TW1 observed that for small adjustments, they felt faster working manually, but for more substantial edits, they preferred using AI assistance.
@@ 3.6.2 How are these challenges addressed?
Iterative refinement of prompts is almost inevitably the programmers’ first recourse when encountering communication challenges with the model. In YT15, when the creator struggled to articulate visual intent for arrow placement, they iterated through multiple prompts with increasingly detailed spoken instructions. Similarly, in YT21, the programmer recognised when initial prompts were too vague and provided more detailed and specific prompts in subsequent attempts, often including numbered lists of requirements and referencing existing code snippets. When the model “did not listen to the system instructions”, TW1 refined the system instructions to emphasise acceptable behaviour; a higher-order strategy than refining individual prompts.
At one point, the programmer in TW1 adopted an unusual prompting strategy where the developer responded to an unsatisfactory result by prompting “Bro, you are losing aura with this. Come on.” before pasting the original prompt and resubmitting. This explicit “vibification” produced more code than the previous turn, though the programmer in TW1 still expressed dissatisfaction with the code quality. This suggests that some vibe coders experiment with social or emotional appeals, albeit perhaps tongue-in-cheek, in prompts. We also observed similar styles of prompting in other videos in our broader corpus. The results of such appeals, as might be expected, are mixed.
Developers strategically switch between different AI models based on their perceived strengths. In YT15, the creator actively switched between OpenAI’s o3-mini and Anthropic’s Claude 3.5 Sonnet models based on their qualitative assessment and the perceived strengths of each for different aspects of the task, using one for animation and another for initial code generation. Similarly, in TW1, the developer discussed where they used models from the Claude family to generate code for a web page, but switched to OpenAI’s GPT 4.5 to rewrite the text because they believed its text quality was an improvement over Claude.
@@ 3.7.1 When and in what ways do experts deploy their expertise in vibe coding?
Expert knowledge also guides problem identification and solution development. In YT22, upon encountering a blank screen, the creator quickly inspects the browser console and identifies a ‘trpc error,’ demonstrating expertise in reading and interpreting error messages. In YT21, the programmer displays expert knowledge of debugging techniques by immediately recognising database issues from error messages, effectively using browser developer tools to diagnose problems, forming hypotheses about root causes, and inserting print statements to trace code execution. YT21 also demonstrates understanding of API endpoints, data structures, and how different parts of the application communicate, allowing them to identify issues with API calls. Similarly, in YT22, the creator identifies that an application is trying to connect to PostgreSQL instead of Supabase and knows exactly how to manually edit environment files to correct the configuration. In YT22, the creator evaluates AI-generated code against expected standards by checking generated React components against documentation, demonstrating and applying their knowledge of UI libraries.
Expertise guides quality control and feature alignment. In YT22, the creator recognizes and decides to remove unwanted AI-generated features that, while not buggy or intrusive, don’t align with the intended design. In TW1, the creator demonstrates understanding of when generated code might incur technical debt, which “lowers developer velocity”, and addresses these issues proactively.
@@ 3.7.2 What kinds of expertise are used in vibe coding?
Beyond “traditional” programming expertise (code quality and maintainability, fault localisation and debugging), the programmers we studied exhibited expertise in two additional domains: AI expertise (understanding models, prompting practices, concepts such as context windows, and limitations of AI code generation), and product management expertise (forming and translating goals into features for the product). Perhaps as a consequence of the blending of prototyping and production (as discussed in Section 3.2.2), vibe coding requires (or at least is greatly facilitated by) a broad skillset spanning traditional programming skills, AI literacy, and product vision.
@@
Moreover, we speculate that these programmers are also developing a sense of what might be called ambient competence. This is a sense of competence wherein the programmer feels capable of tackling tasks they wouldn’t have considered before, because they have access to an AI system that may be able to accomplish the task with very little effort invested, regardless of whether they themselves have the skills to tackle them manually (which they very well may). This can be thought of as the competence that arises as the result of the active fulfilment of the “awareness of the possible” (Sarkar, 2023c, ; Sarkar, 2023b, ). This could lead to a shifting locus of agency, where the programmer’s confidence and the scope and complexity of projects they undertake are increasingly defined by their perceived ability to prompt and guide a model, rather than their own direct coding abilities.
@@ 3.7.3 When do practitioners transition from vibe coding to manual work?
For straightforward edits where the overhead of prompting exceeds the effort of making the change directly, developers often opt for manual work. In YT21, the programmer attempted an inline edit with AI but then decided to perform the simple one-line edit manually after determining that AI assistance was unnecessary for such a minor change. In TW1, the developer articulates their decision-making process about when to work manually versus using AI assistance. In one instance, they switch to manual work when they believed it would be a more efficient workflow, noting that “IntelliJ’s autocomplete and search is still faster than LLMs”. This leads them to “race the AI” to find code locations, sometimes stopping AI generation mid-process when they locate what they need faster.
@@
Another common transition point across videos is when AI generates a mostly-complete solution that requires refinement. This can be viewed as a special case of transitioning to manual work for efficiency. For example, TW1 accepted code when the AI “got close” to their goal and then manually edited the generated code to be more inline with their preferences for code. TW1 describes their viewpoint that LLMs should be used to “get close to what you want” but that expecting perfect outputs is “a mistake” – instead, their optimal workflow involved making post-generation adjustments to generated code, and reuse of generated code in similar features by quickly modifying code to other contexts rather than issuing refined prompts.
Similarly, in YT22, the creator performs manual edits to remove unwanted UI elements, delete redundant buttons, and make small edits in the code to better align with their vision. For instance, the creator manually removes functionality for uploading job descriptions via URL to focus solely on text-based inputs, representing a conscious design decision that required direct code manipulation.
@@
Debugging is as a common trigger for manual intervention across multiple videos. In YT21, the programmer switches to manual work almost immediately when encountering errors, using traditional debugging techniques like browser developer tools, console inspection, and hypothesis formation. In one instance, they hypothesised a solution based on an error message and implemented a fix without AI assistance. Throughout, YT21 frequently engages in manual code review and rejection of AI suggestions based on their understanding of correctness and project requirements.
YT22 shows similar patterns, with the developer manually editing environment files to fix database connection errors and directly intervening to correct runtime issues. YT22 also manually corrects a misalignment in AI-generated documentation (changing "Next.js 14" to "Next.js 15") and deletes redundant steps in implementation plans. These examples suggest that when errors occur, developers often rely on their expertise and traditional debugging approaches rather than AI tools.
@@
However, YT15 presents a contrasting case where the developer states explicitly that they are “not going to try and fix it on his own” when encountering an error, citing adherence to the Karpathy canon. The creator in YT15 deliberately avoids manual debugging and relies almost exclusively on re-prompting for changes throughout their workflow. Thus, the videos do exhibit individual differences in approaches to manual work.